OpenStudy (anonymous):

Let a,b,c be positive real no.s such that a^3+b^3=c^3.Prove that ..

OpenStudy (anonymous):

wait a moment

OpenStudy (amistre64):

i dont think you can prove it beyond R^2

OpenStudy (zarkon):

they are real numbers not integers

OpenStudy (amistre64):

volume behaves differently than area

OpenStudy (anonymous):

\[a ^{2}+b ^{2}-c^{2}>6(c-a)(c-b)\]

OpenStudy (across):

Have you read about Fermat's Last Theorem? ;P

OpenStudy (zarkon):

doesn't apply here

OpenStudy (amistre64):

im waiting for the movie :)

OpenStudy (across):

I know (we're talking about real numbers), but it reminded me of it. ;P Me too!

OpenStudy (zarkon):

if you want to know about Fermat's Last Theorem read Fermat's Enigma.

OpenStudy (anonymous):

u saw that google doodle about fermat's theorem

OpenStudy (anonymous):

When is the movie coming out amistre ? I hope they didn't make it a vampire movie (too many of them recentlu)

OpenStudy (across):

If \[a^3+b^3=c^3,\]then \[a^2+b^2-c^2 \gt 6(c-a)(c-b).\]You can prove this by way of a counter-example: Let \[a=0.\]We then see that \[b^3=c^3,\]\[b=c.\]Consequently, \[b^2-c^2 \gt 6(c-a)(c-b),\]\[b^2-b^2 \gt 6c(c-b),\]\[0 \gt 6b(b-b),\]\[0 \gt 0,\]and we have encountered a contradiction. Therefore, the conclusion is false. QED

OpenStudy (zarkon):

@ across "Let a,b,c be positive real "